An Earth satellite is orbiting at a distance from the Earth's surface equal to one Earth radius.  At this location, the acceleration due to gravity is what factor times the value of g at the Earth's surface?
Answer #1
Graviatational force on a mass m at r = R + R = 2R from thecenter of the earth,
         F' =mg' = G.M.m/4R2
Graviatational force on the same mass m at r = R (on thesurface of the earth).
        F = mg =G.M.m/R2
Dividing,
          g'/g =1/4 = 0.25
         g' = 0.25g